Evaluate ∭E  (𝑥z−y3)dV where E = { (x,y,z) -1 ≤ x ≤ 1,  0 ≤ y ≤ 2,  0 ≤ z ≤ 1 }.


1
Expert's answer
2022-01-20T15:31:05-0500

Solution:

"\begin{aligned}
&E=\left\{\int_{E}(x, y, z) /-1 \leq x \leq 1,0 \leq y \leq 2,0 \leq z \leq 1\right\} \\
&\iiint_{0}\left(x z-y^{3}\right) d V=\int_{0}^{1} \int_{0}^{2} \int_{-1}^{1}\left(x z-y^{3}\right) d x d y d z \\
&=\int_{0}^{1} \int_{0}^{2}\left(\frac{x^{2}}{2} z-y^{3} x\right)_{x=-1}^{x=1} d y d z \\
&=\int_{0}^{1} \int_{0}^{2}\left[\left(\frac{1}{2}-\frac{1}{2}\right) z-y^{3}(1+1)\right] d y d z \\
&=\int_{0}^{1} \int_{0}^{2}\left(-2 y^{3}\right) d y d z \\
&\text { i.e. } \left.\iiint_{E}\left(x z-y^{3}\right) d V=\int_{0}^{1}-2 \frac{y^{4}}{4}\right]_{y=0}^{2} d z \\
&=\int_{0}^{1}-\frac{2}{4}(16-0) d z \\
&\left.=-8 \int_{0}^{1} d z=-8 z\right]_{0}^{1} \\
&=-8(1-0) \\
&=-8
\end{aligned}"

"\begin{aligned}
&\text { Now } \iiint_{E}\left(x z-y^{3}\right) d V=\int_{0}^{2} \int_{1}^{1} \int_{0}^{1}\left(x z-y^{3}\right) d z d x d y\\
&=\int_{0}^{2} \int_{-1}^{1}\left[x \frac{z^{2}}{2}-y^{3} z\right]_{z=0}^{z=1} d x d y\\
&=\int_{0}^{2} \int_{-1}^{1}\left(x \frac{1}{2}-y^{3}\right) \cdot d x d y\\
&=\int_{0}^{2}\left[\frac{x^{2}}{4}-y^{3} x\right]_{x=-1}^{x=1} d y\\
&=\int_{0}^{2}\left(\frac{1}{4}(1-1)-y^{3}(1+1)\right) d y\\
&\left.=\int_{0}^{2}(-2) y^{3} d y=(-2) \frac{y^{4}}{4}\right]_{0}^{2}\\
&=\frac{-2}{4}(16-0)=-8\\
&\text { And } \iiint_{E}\left(x z-y^{3}\right) d V=\int_{-1}^{1} \int_{0}^{1} \int_{0}^{2}\left(x z-y^{3}\right) d y d z d x\\
&=\int_{-1}^{1} \int_{0}^{1}\left[x z y-\frac{y^{4}}{4}\right]_{y=0}^{y=2} d z d x\\
&=\int_{-1}^{1} \int_{0}^{1}\left[x z 2-\frac{16}{4}\right] d z d x\\
&=\int_{-1}^{1}\left[2 x \cdot \frac{z^{2}}{2}-4 z\right]_{z=0}^{z=1} d x
\end{aligned}"

"\text { i.e. }"

"\begin{aligned}
& \iiint_{E}\left(x z-y^{3}\right) d V=\int_{-1}^{1}[x(1)-4(1-0)] d x \\
=& \int_{-1}^{1}(x-4) d x \\
=&\left[\frac{x^{2}}{2}-4 x\right]_{x=-1}^{x=1} \\
=& \frac{1}{2}(1-1)-4(1+1) \\
=& 0-4(2) \\
=&-8
\end{aligned}"